OpenStudy (anonymous):

Medal! The equation below shows the area of a trapezoid, A, with a height of 9 cm, and one base 35 cm. A = 9 over 2(b + 35) Which of the following formulas correctly solves for the other base, b? b = 2A over 9 + 35 b = 2 multiplied by A over 9 - 35 b = 2 multiplied by A plus 35, all over 9 b = 2 multiplied by A minus 35, all over 9

OpenStudy (***[isuru]***):

hi, \[A = \frac{ 9 }{ 2 }(b + 35)\] \[(b + 35 )= \frac{ 2A }{ 9 }\] \[b = \frac{ 2A }{ 9 } -35\] Yep!! that;s it and now I think ya can guess the answer!!!'' Hope this will help ya out!!!